OpenStudy (anonymous):

evaluate f(g(0)) and g(f(1)) where f(x)=sgrt(x^2-1) and g(x)=(x^2+1)/(x^2-1)

OpenStudy (anonymous):

i think i got f(g(x)) to be sqrt((x^2+1)/(x^2-1)-1) and g(f(x)) to be sqrt(((x^2+1)/(x^2-1))^2-1)

hartnn (hartnn):

not required.. can u find g(0) =?

OpenStudy (anonymous):

well i think g(0) would be -1

hartnn (hartnn):

yup, thats correct, now find f(-1)

OpenStudy (anonymous):

which if you plug -1 into f(x) would be 0?

hartnn (hartnn):

yup, so f(g(0)) = 0 got this?

OpenStudy (anonymous):

so g(f(1)) would be -1?

hartnn (hartnn):

yup

OpenStudy (anonymous):

awesome thanks!

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

by chance want to help with another problem?

hartnn (hartnn):

by every chance.....

OpenStudy (anonymous):

2^(1+cosx)=4

OpenStudy (anonymous):

i think the next step is log(base 2)1+cosx=log(base 2)4

hartnn (hartnn):

write 4 as 2^2

hartnn (hartnn):

no need to take log 2^(1+cos x) = 2^2 so 1+cos x =2 got this?

OpenStudy (anonymous):

soo cosx=1?

hartnn (hartnn):

yup

OpenStudy (anonymous):

i think im thinking too hard ha. thanks!

hartnn (hartnn):

welcome :)